### v1

# Illumination number of 3-dimensional cap bodies

Abstract.
Despite recent progress on the Illumination conjecture, it remains open in general, as well as for specific classes of bodies. Bezdek, Ivanov, and Strachan showed that the conjecture holds for symmetric cap bodies in sufficiently high dimensions. Further, Ivanov and Strachan calculated the illumination number for the class of 3-dimensional centrally symmetric cap bodies to be 6.
In this paper, we show that even the broader class of all 3-dimensional cap bodies has the same illumination number 6, in particular, the illumination conjecture holds for this class. The illuminating directions can be taken to be vertices of a regular tetrahedron, together with two special directions depending on the body. The proof is based on probabilistic arguments and integer linear programming.
Key words and phrases:
Illumination problem, Hadwiger’s covering conjecture, cap body, random rotations, integer linear programming.
2020 Mathematics Subject Classification:
Primary 52A15; Secondary 52A40, 52C17, 90C05
The second author was supported in part by NSERC of Canada Discovery Grant RGPIN-2020-05357.
The third author was supported by NSERC of Canada Discovery Grant RGPIN-2020-05357.
1. Introduction
Hadwiger [1] asked the following question in 1957: For $n\geq 3$, what is the smallest number $N(n)$ such that every $n$-dimensional convex body can be covered by the union at most $N(n)$ of translates of the body’s interior? The earliest formulation of the question dates back to 1955 [2], where Levi proved that $N(2)=4$. After Hadwiger and Levi, the question was restated by Gohberg and Markus [3] (independently without knowing about the work of Levi and Hadwiger) in terms of covering by homothetic copies. For more details about the history of the problem and an extensive list of partial results, see, e.g., [BK].
The example of the $n$-dimensional cube yields $N(n)\geq 2^{n}$. It is widely believed that $N(n)=2^{n}$ for all $n\geq 2$.
Conjecture 1.
Any $n$-dimensional convex body can be covered by $2^{n}$ or fewer smaller positive homothetic copies of itself, $n\geq 3$. Furthermore, $2^{n}$ homothetic copies are required only if the body is an affine copy of the $n$-cube.
Boltyanski [B] showed that the conjecture is equivalent to a certain illumination problem. For a convex body $K$, a direction (unit vector) $v$ illuminates a point $x$ on the boundary $\partial K$ of $K$, if the ray $\{x+vt\;:\;t\geq 0\}$ has nonempty intersection with interior of $K$. The set of directions $\{v_{i}\}_{i=1}^{k}$ is said to illuminate $K$ if every point of $\partial K$ is illuminated by some $v_{i}$. The illumination number $I(K)$ of $K$ is the smallest $k$ for which $K$ can be illuminated by $k$ directions. Boltyanski [B] showed that  1 is equivalent to the following: the illumination number of a convex body in $\mathbb{E}{}^{n}$ does not exceed $2^{n}$, with equality only for affine copies of the $n$-cube.
In literature, 1 has been referred to as the Hadwiger Covering Conjecture, the Levi-Hadwiger(-Gohberg-Markus) Conjecture, and Hadwiger-Boltyanski Illumination Conjecture.
As of today, the conjecture has only been solved for the $2$-dimensional case, and it has been notoriously hard to crack even for the next smallest dimension, namely, $\mathbb{E}^{3}$. Several people have worked on the $3$-dimensional case. Papadoperakis [6] proved that the illumination number of any convex body in ${\mathbb{E}{}}^{3}$ is at most 16, and Prymak [Pr] improved the bound to 14. This result is far from the expected illumination number $8$, but it is the best known upper bound on illumination number in $\mathbb{E}{}^{3}$.
The conjecture has been verified for certain classes of bodies in $\mathbb{E}{}^{3}$. Bezdek and Kiss [BKiss] showed that the illumination number of almost smooth convex bodies and the illumination number of convex bodies of constant width is at most 6. The result about bodies of constant width has been proven independently by several people (refer to [BLNP] for more details). Martini [M] proved that the illumination number of every zonotope in $\mathbb{E}^{3}$, which is not a parallelotope, is at most 6. Lassak [11] proved that if the convex body is centrally symmetric, then the conjecture holds in $\mathbb{E}{}^{3}$. Furthermore, Dekster [12] proved (without treatment of the equality case) that the conjecture holds in $\mathbb{E}{}^{3}$ if the convex body is symmetric about a plane.
A convex body $K$ is called a cap body if and only if $K=\bigcup_{i=1}^{m}\text{conv}(\{x_{i}\}\cup\mathbb{B}^{n})$ for some points $x_{i}\in\mathbb{E}{}^{n}\setminus\mathbb{B}^{n}$, where $\mathbb{B}^{n}$ is the unit sphere. Bezdek, Ivanov and Strachan [BIS] proved the conjecture for centrally symmetric cap bodies in sufficiently large dimensions. In particular, they proved that any $n$-dimensional centrally symmetric cap body can be illuminated by $<2^{n}$ directions in Euclidean $n$-space for $n=3,$ $4,$ $9$ and $n\geq 19$. Moreover, Ivanov and Strachan [IS] proved that any $3$-dimensional centrally symmetric cap body can be illuminated by $6$ directions. This bound cannot be improved: the convex hull of the unit ball and an appropriately scaled regular octahedron is a centrally symmetric cap body with illumination number exactly $6$. In this work, we use a completely different method and show that for any convex body from the much larger class of not necessarily symmetric $3$-dimensional cap bodies, the illumination number still does not exceed $6$. This is sharp by the same example.
Theorem 2.
For any cap body $K\subset\mathbb{E}{}^{3}$, $I(K)\leq 6$.
In particular, Theorem 2 verifies the conjecture for all cap bodies in $\mathbb{E}^{3}$. The illuminating directions can be taken to be vertices of a regular tetrahedron, together with two special directions depending on the body. The existence of such a tetrahedron is shown probabilistically, with an estimate on the expected number of un-illuminated caps (which turns out to be $\leq 2$) carried over using integer linear programming, which allows delegating analysis of cases of the quantities of caps of various sizes to the computer.
Remark 3.
We shall address the illumination problem for cap bodies in higher dimensions in a forthcoming paper.
2. Preliminaries
[…]
Let $\text{conv}(X)$ be the convex hull of the set $X$. $K$ is a convex body in $\mathbb{E}{}^{n}$ if it is convex compact set with non-empty interior. A convex body $K$ is called a cap body if and only if
$K=\bigcup_{i=1}^{m}\text{conv}(\{x_{i}\}\cup\mathbb{B}^{n})$
for some points $x_{i}\in\mathbb{E}{}^{n}\setminus\mathbb{B}^{n}$, $1\leq i\leq m$, which are called vertices of $K$. For a given vertex $x_{i}$, the corresponding base cap (or simply cap) is defined to be the set
$S_{i}:=\text{conv}(\{x_{i}\}\cup\mathbb{B}^{n})\cap\mathbb{S}^{n-1}.$
Note that $S_{i}=C[\widehat{x}_{i},\varphi_{i}]$, where $\widehat{x}_{i}:=\frac{x_{i}}{\|x_{i}\|}$ is the centre of the cap, and $\varphi_{i}=\arccos{}\frac{1}{\|x_{i}\|}$ is the radius of the cap. It is an easy observation that the base caps always have acute radius, i.e., $\varphi_{i}<\pi/2$ for $1\leq i\leq m$. Furthermore, observe that the convexity of $K$ implies that $C(\widehat{x}_{\alpha},\varphi_{\alpha})\cap C(\widehat{x}_{\beta},\varphi_{\beta})=\emptyset$ for any distinct $x_{\alpha},x_{\beta}\in\{x_{i}\}_{i=1}^{m}$. One can construct a cap body from any given system of mutually non-overlapping open caps of radius $<\tfrac{\pi}{2}$ on the sphere. In particular, the vertex $x_{i}$ can be expressed using $C(\widehat{x}_{i},\varphi_{i})$ as follows:
$x_{i}=\frac{\widehat{x}_{i}}{\cos{\varphi_{i}}}.$
We use the following proposition from [BIS, ABPR]:
Proposition 4.
A cap body $K$ with vertices $\{x_{i}\}_{i=1}^{m}$ is illuminated by the directions $\{v_{j}\}^{k}_{j=1}\subset\mathbb{S}^{n-1}$ if:
  1. i)
    $C\left(-\widehat{x}_{i},\frac{\pi}{2}-\varphi_{i}\right)\cap\{v_{j}\}^{k}_{j=1}\neq\emptyset$ for each $i$, $1\leq i\leq m$,
  2. ii)
    positive hull of $\{v_{i}\}^{k}_{i=1}$ is $\mathbb{E}{}^{n}$, i.e. for all $x\in\mathbb{E}{}^{n}$ there are positive $c_{1},\ldots,c_{k}$ such that $x=c_{1}v_{1}+\cdots+c_{k}v_{k}$.
We will say that the cap $S_{i}$ is illuminated by $v_{j}$ if and only if $v_{j}\in C\left(-\widehat{x}_{i},\frac{\pi}{2}-\varphi_{i}\right)$.
3. Proof of Theorem 2
First, let us prove the existence of a cap body $K_{0}$ such that $I(K_{0})=6$. This was already established by Ivanov and Stranchan [IS], but we include the example here for the sake of completeness. Define $S:=\{\pm(1,0,0),\pm(0,1,0),\pm(0,0,1)\}$, and construct $K_{0}$ by taking $6$ caps (open caps) with radius $\frac{\pi}{4}$ and centers at each element of $S$. As per Proposition 4, we need to find a set $V=\{v_{i}\}_{i=1}^{k}$ such that $C\left(-x,\frac{\pi}{4}\right)\cap V\neq\emptyset$ for each $x\in S$, and positive convex hull of $\{v_{i}\}^{k}_{i=1}$ is $\mathbb{E}^{3}$. Note that the caps $C\left(-x,\frac{\pi}{4}\right),\;x\in S$ do not intersect, so each cap requires at least one direction. So, $|V|\geq 6$. It is easy to see that $V=S$ illuminates the body, as positive hull of $S$ is $\mathbb{E}^{3}$. This proves that $I(K_{0})=6$.
Now, suppose $K$ is a cap body with vertices $x_{1},x_{2},\dots,x_{m}$, and we assume that sizes of caps are ordered $\varphi_{1}\geq\varphi_{2}\geq\ldots\geq\varphi_{m}$. We will show $I(K)\leq 6$.
Let $L$ be a set of vertices of a regular tetrahedron inscribed into the unit sphere. It is routine work to show that $L$ satisfies Proposition 4 (ii), and thus we only need to check Proposition 4 (i) to complete the proof.
For every $\theta\in(0,\frac{\pi}{2}]$, define $C_{\theta}:=\bigcup_{l\in L}C[l,\theta]$. It is well-known and is a simple computation that $C_{\theta}=\mathbb{S}^{2}$ when $\theta\geq\arccos\left({\frac{1}{3}}\right)$. Therefore, by Proposition 4 (i), any vertex $x_{j}$ of $K$ with $\varphi_{j}<\frac{\pi}{2}-\arccos\left({\frac{1}{3}}\right)$ is illuminated by one of the directions from any rotation of $L$. Thus, in what follows, we assume that $\varphi_{m}\geq\frac{\pi}{2}-\arccos\frac{1}{3}$.
Let $\sigma$ be the probabilistic spherical measure on $\mathbb{S}^{2}$. To find a suitable random rotation that illuminates most of the other caps (with radius at least $\psi_{0}$), we need to know the value of $\sigma(C_{\theta})$.
Lemma 5.
We have:
(1) $\sigma(C_{\theta})=$ $2(1-\cos\theta),$ $0<\theta\leq\tfrac{1}{2}\arccos(-\tfrac{1}{3})$,
(2) $\sigma(C_{\theta})=$ $2(1-\cos\theta)-6A_{\theta},$ $\tfrac{1}{2}\arccos(-\tfrac{1}{3})\leq\theta<\arccos\tfrac{1}{3}$,
(3) $\sigma(C_{\theta})=$ $1,$ $\arccos\tfrac{1}{3}\leq\theta\leq\tfrac{\pi}{2}$,
where
$A_{\theta}=\frac{1}{2\pi}\left(-\arccos\left(\frac{-\frac{1}{3}-\cos^{2}\theta}{\sin^{2}\theta}\right)-2\arccos\left(\sqrt{2}\cot\theta\right)\cos\theta+\pi\right).$
Proof.
Recall that $C_{\theta}=\mathbb{S}^{2}$ for $\theta>\frac{\pi}{2}-\psi_{0}$, so (3) readily follows. (1) is straightforward from the standard $\sigma(C[x,\theta])=\frac{1-\cos\theta}{2}$ and the fact that the four congruent caps in $C_{\theta}$ do not overlap in this case. For (2), we observe that there will be six congruent “lunes” which are intersections of two caps from $C_{\theta}$. The measure $A_{\theta}$ of each lune can be computed using, for example [TV]*Eqs. (2)–(4). ∎
Next, we will upper bound the expected number of caps not illuminated by a random rotation of $L$. This will be done utilizing integer programming. We need to consider various cases depending on the number of caps of different sizes. Set $a_{0}:=\tfrac{19\pi}{180}<\tfrac{\pi}{2}-\arccos\tfrac{1}{3}\leq\varphi_{m}$. For a suitable positive integer $t$ that will be selected later, we define a discretization array $a=[a_{0},a_{1},\dots,a_{t}]$, where $a_{t}:=\frac{\pi}{2}$ and $a_{i}=a_{0}+i\frac{a_{t}-a_{0}}{t}$ for $11\leq i\leq t-$ 1 are equidistant on $[a_{0},a_{t}]$. We have $\varphi_{j}\in(a_{0},a_{t}]$ for any $j$. Let $n_{i}$, $10\leq i\leq t-$ 1, denote the number of indices $j$ such that $\varphi_{j}\in(a_{i},a_{i+1}]$. Note that by Proposition 4 (i) a vertex $x_{j}$ with $\varphi_{j}\in(a_{i},a_{i+1}]$ is illuminated by a random rotation $L^{\prime}$ of $L$, provided at least one of the points of $L^{\prime}$ is within the geodesic distance $\tfrac{\pi}{2}-\varphi_{j}$ of $\widehat{x_{j}}$. Therefore, the probability that $x_{j}$ is not illuminated is $1-\sigma\left(C_{\tfrac{\pi}{2}-\varphi_{j}}\right)\leq 1-\sigma\left(C_{\tfrac{\pi}{2}-a_{i+1}}\right)$. Then overall, the expected number of caps which are not illuminated does not exceed
(4) $\sum_{i=0}^{t-1}n_{i}\left(1-\sigma\left(C_{\tfrac{\pi}{2}-a_{i+1}}\right)\right),$
which will be our target function in the integer programming problem.
Due to convexity of $K$, we know that the caps $C(\widehat{x}_{j},\varphi_{j})$ do not overlap, so the total measure of these caps is at most $1$. In terms of $n_{i}$, this provides the following constraint:
(5) $\sum_{i=0}^{t-1}n_{i}\frac{1-\cos a_{i}}{2}\leq 1.$
To obtain another constraint, we need the following lemma which states that one cannot pack five caps of radius $>\tfrac{\pi}{4}$ on $\mathbb{S}^{2}$. This is likely not new, but we include the short proof anyway for completeness.
Lemma 6.
If $m\geq 5$, then $\varphi_{5}\leq\frac{\pi}{4}$.
Proof.
Assume to the contrary that $\varphi_{5}>\frac{\pi}{4}$, hence $\varphi_{j}>\frac{\pi}{4}$, $1\leq j\leq 5$. As $K$ is a cap body, all base caps $C(\widehat{x}_{j},\varphi_{j})$, $1\leq j\leq 5$, are disjoint, so $\theta(x_{i},x_{j})\geq\varphi_{i}+\varphi_{j}>\frac{\pi}{2},$ and so $\langle\widehat{x}_{i},\widehat{x}_{j}\rangle<0$, $1\leq i<j\leq 5$. The points $\widehat{x}_{1},\widehat{x}_{2},\dots,\widehat{x}_{5}$ are affinely dependent, so there are $c_{1}$, $c_{2}$, $c_{3}$, $c_{4}$, $c_{5}$ (not all zero) such that
$c_{1}\widehat{x}_{1}+...+c_{5}\widehat{x}_{5}=0$
and $\sum_{i=1}^{5}c_{i}=0$. Suppose that $I_{+}=\{c_{i}:c_{i}\geq 0\}$ and $I_{-}=\{c_{j}:c_{j}<0\}$, now
$\sum_{c_{i}\in I_{+}}c_{i}\widehat{x}_{i}=\sum_{c_{j}\in I_{-}}-c_{j}\widehat{x}_{j}.$
Finally, taking the dot product with $\sum_{c_{i}\in I_{+}}c_{i}\widehat{x}_{i}$ on both sides we get
$0\leq\left\|\sum_{c_{i}\in I_{+}}c_{i}\widehat{x}_{i}\right\|^{2}=\sum_{c_{i}\in I_{+}}c_{i}\widehat{x}_{i}\sum_{c_{j}\in I_{-}}-c_{j}\widehat{x}_{j}=\sum c_{i}(-c_{j})\langle\widehat{x}_{i},\widehat{x}_{j}\rangle<0,$
which is the desired contradiction. ∎
In terms of $n_{i}$, Lemma 6 implies that
(6) $\sum_{0\leq i<t\,:\,a_{i}\geq\tfrac{\pi}{4}}n_{i}\leq 4.$
Let $M_{t}$ denote the solution of the integer linear programming problem
maximize (4) subject to (5) and (6)
with non-negative integer variables $n_{i}$, $10\leq i\leq t-$ 1. If for some $t$ we get $M_{t}<3$, then there exists a random rotation of $L$ such that the expected number of caps not illuminated by the rotation is at most $\lfloor M_{t}\rfloor\leq 2$. Assigning a direction for each of these caps, we obtain an illuminating system with at most $6$ directions.
For any given $t$, the computation of $M_{t}$ is a standard integer linear program for which many solvers are available. We are interested in an upper bound on $M_{t}$. Thus, to avoid any numerical errors, we can fix a positive integer denominator $D$ and round up the coefficients (computed symbolically) in (4), and round down the coefficients in (5) to the nearest number from $\frac{1}{D}\mathbb{Q}$. The obtained modified integer linear program will have rational coefficients and can be solved precisely without numerical errors. The corresponding SageMath [sagemath] script is given in the appendix. With $D=3000$ and $t=250$, the solution to the modified problem is $\frac{2999}{1000}$, and thus $M_{250}\leq 2.999$ implying the desired $\lfloor M_{250}\rfloor\leq 2$ and completing the proof.
Remark 7.
The script takes less than 2 hours to complete on a modern personal computer. If we switch to computations with floating point arithmetics, the approximate solution can be obtained significantly faster (under 1 second for the same $t=250$). Such computations with much larger $t$ suggest that $\limsup_{t\to\infty}M_{t}<2.97$. We emphasize that the bound $M_{250}\leq 2.999$ in our proof was obtained using symbolic computations only. One can also get the required bound using fewer intervals than $250$ by partitioning $[a_{0},\frac{\pi}{2}]$ in a non-uniform manner. We opted to use the uniform partition and simplicity of the exposition at the cost of a slight increase in computation time.
Remark 8.
If we aim to show that $I(K)\leq 7$, then a human verifiable proof is possible. The proof would require adding one geometric argument and would make the optimization problem reducible to under a dozen cases.
{bibsection}
Appendix
{python}
a0 = 19*pi/180 at = pi/2 t = 250 a = [a0+i*(at-a0)/t for i in range(t+1)]
D=3000 #denominator for rationalization, rounding functions def up(x): return(ceil(x*D)/D) def down(x): return(floor(x*D)/D)
p = MixedIntegerLinearProgram(solver=”PPL”) #exact computations for programs with rational coefficients v = p.new_variable(integer=True, nonnegative=True)
#v[i] is the number of base caps with angle between a[i] and a[i+1]
def mu(x): th = pi/2-x #computing the formula in Lemma 5 for th if th¡arccos(-1/3)/2: return 2*(1-cos(th)) if th¡arccos(1/3): return 2*(1-cos(th))-3/pi*(-arccos((-1/3-(cos(th))^2)/(sin(th))^2)-2*arccos(sqrt(2)*cot(th))*cos(th)+pi) return 1
p.set_objective(sum(v[i]*(up(1-mu(a[i+1]))) for i in range(t))) #this upper bounds (rounding up) the target function from (4) p.add_constraint(sum(v[i]*(down(1-cos(a[i]))) for i in range(t))¡=2) #constraint (5), rounding down to include all feasible solutions of the problem in the paper p.add_constraint(sum(v[i] for i in range(t) if a[i]¿=pi/4)¡=4) #constraint (6)
answer=p.solve() print(answer,floor(answer))